# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v1

@@ 3.1 Preliminary
Let $\boldsymbol{\mu}=\begin{bmatrix}\mathbb{E}[f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x})],&\mathbb{E}[f_{\boldsymbol{\theta}_{2}}(\boldsymbol{x})],&\dots&\mathbb{E}[f_{\boldsymbol{\theta}_{N}}(\boldsymbol{x})]\end{bmatrix}^{\top}$ be the vector of expected returns of the base learners, $\boldsymbol{\Sigma}$ be the covariance matrix where $\Sigma_{ij}=\text{Cov}(f_{\boldsymbol{\theta}_{i}}(\boldsymbol{x}),f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}))$, and $R_{T}$ be the target return for the portfolio. Let $\boldsymbol{\hat{w}}$ be an estimator of $\boldsymbol{w}$ based on an i.i.d. sample from a distribution $\mathcal{X}$. The ensemble prediction error can be decomposed into bias and variance components. In this case, the bias is minimized subject to the portfolio return being above the target value, and the variance is minimized. The optimization problem can be formulated as follows.
@@
where $\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-R_{T}\right)^{2}$ represents the bias of the expected return of the portfolio compared to the target return, and $\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ represents the variance of the portfolio. $\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq R_{T}$ ensures that the expected return of the portfolio is above the target return. This setup allows to find the optimal weights $\boldsymbol{\hat{w}}$ that minimize the total prediction error of the portfolio, considering both bias and variance, while ensuring that the expected return of the portfolio is above the target value. The expression in (1) generalizes the standard MSE (when $A=I$), reflecting the bias-variance trade-off (Cai et al., 2024):
$MSE_{A}(\boldsymbol{\hat{w}})=\mathcal{L}_{A}(\mathbb{E}[\boldsymbol{\hat{w}}]-\boldsymbol{w})+\text{tr}(A\text{Var}[\boldsymbol{\hat{w}}])$ (2)
where the first term vanishes for an unbiased estimator, and the second term vanishes if the estimator has zero variance. Now, consider $\boldsymbol{w}$ as the solution to a decision problem:
@@
where the expectation is taken under $D$. For a given $\boldsymbol{\hat{w}}$ and matrix $A$, establish a relationship between $MSE_{A}(\boldsymbol{\hat{w}})$ and $D_{P}(\boldsymbol{\hat{w}})$, expressed as a function $f$:
$MSE_{A}(\boldsymbol{\hat{w}})=f(D_{P}(\boldsymbol{\hat{w}}))$ (5)
For a parameterized estimator $\boldsymbol{\hat{w}}(\delta)$, compare $\delta_{MSE}=\arg\min_{\delta}MSE_{A}(\boldsymbol{\hat{w}}(\delta))$ with $\delta_{D}=\arg\min_{\delta}D_{P}(\boldsymbol{\hat{w}}(\delta))$ and assess the decision loss increase:
@@ 3.1.1 Diversity in Ensemble Learning
where $\overset{\circ}{\boldsymbol{R}}_{j}=\arg\min\mathbb{E}_{D}[\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{R}_{j})]$ and the combiner is
$\boldsymbol{\hat{R}_{T}}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{R}_{j})$, equivalent to the equal-weighted-portfolio ($1/N$-portfolio rule (DeMiguel et al., 2009)). This decomposition separates the expected ensemble risk into: the noise, which represents the irreducible error inherent in the problem; the average bias of the portfolio returns predictions $\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R_{T}},\overset{\circ}{\boldsymbol{R}}_{j})$; the average variance of the individual models’ predictions in the portfolio ensemble $\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}[\mathcal{L}(\overset{\circ}{\boldsymbol{R}}_{j},\boldsymbol{R}_{j})]$; the expected ambiguity term, considered as the ensemble diversity, which has the opposite sign to the other terms, $\mathbb{E}\left[\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{\hat{R}_{T}},\boldsymbol{R}_{j})\right]$, capturing the disagreement among the portfolio ensemble members in their predictions.
@@
3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
In this portfolio structured ensemble model framework, the goal is to make portfolio allocation decisions at time $t$ for an investment horizon $\Delta t=T-t$. The daily time-series returns of a portfolio for the period $\Delta t$ are denoted as $\boldsymbol{R_{T}}=\{R_{Ti}\}_{i=t}^{T}$. To approximate the expected return, a structured ensemble model predicts the portfolio time-series returns as $\boldsymbol{\hat{R}_{T}}$. This model consists of $M$ individual predictors, $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x_{j}})\}_{j=1}^{M}$, parameterized by $\Theta=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, where each predictor focuses on forecasting the return of the $j$-th component of the portfolio $\{\boldsymbol{R}_{j}\}_{j=1}^{M}$. The predictors are trained using stochastic gradient descent (SGD). During each training iteration $i$, for a given input $\{\boldsymbol{x}_{ij}\}_{i=1,j=1}^{N,M}$ and its corresponding target $R_{ij}$, the parameters of the $j$-th predictor are updated as:
where $\mathcal{F}$ is a set of specific functions of the ensemble model, $\boldsymbol{w}=\{w_{j}\}_{j=1}^{M}$ represents the parameters of the ensemble, which, when appropriately scaled, are equivalent to the portfolio weights, and $\boldsymbol{\hat{R}_{T}}$ represents the predicted time series returns for the period $\Delta t$. By incorporating the diversity parameter $\varepsilon$, the ensemble achieves a balance between prediction accuracy and diversity, enhancing its generalization capabilities in portfolio forecasting tasks. Interestingly, this diversity aligns with the concept of portfolio diversification, as originally proposed in Modern Portfolio Theory (Markowitz, 1952).
Figure 1: Structured data set formation, including diversity ($\varepsilon$), and analytical ensemble training of the s-RBFN model, described in Section 3.3 (Dominguez et al., 2025). $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{i})$ is the $j^{th}$ shooter (predictor), $\hat{R}_{ij}$ is the $i^{th}$ shoot of $j^{th}$ shooter and $\boldsymbol{\hat{R}}=\boldsymbol{\hat{R}}(\varepsilon)$ is the full set of shoots after training, $\boldsymbol{R_{T}}$ is the portfolio target, and $\boldsymbol{w}$ is the s-RBFN model parameters optimized by least-squares.
@@ 3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
The optimal s-RBFN parameters can now be obtained by least-squares with the regularization parameter $\lambda_{s}$ and target portfolio $\boldsymbol{y}$:
$\boldsymbol{w}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{y}$ (13)
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, including ensuring they sum to one and remain within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\mathbf{w}=1$ and $\mathbf{0}\leq\mathbf{w}\leq\mathbf{1}$.
@@
$\boldsymbol{w}=\boldsymbol{P}\boldsymbol{R_{T}}=\boldsymbol{P}\boldsymbol{w}^{T}\boldsymbol{x}$ (14)
with $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}$, and $\boldsymbol{w}$ assumed to be optimal. In this case, the shape of the constituents’ hypothesis space and the degree of diversification in the structured data set of predictions have already been incorporated into $\boldsymbol{P}$. To ensure that this information is utilized in the optimization of the s-RBFN parameters (portfolio ensemble model weights from (13)), regardless of the terms $\boldsymbol{w}^{T}\boldsymbol{x}$ in (14) for the target portfolio, normalization is applied so that an equally weighted target portfolio $\boldsymbol{R_{T}}$ is assumed, with weights $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the s-RBFN parameters obtained during training are:
$\boldsymbol{\hat{w}}\approx\boldsymbol{w}=\boldsymbol{P}\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ (15)
The Centroid Combiner Rule, derived from the Bias-Variance-Diversity decomposition in (8) and adapted to this problem,
$\boldsymbol{\hat{R}_{T}}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{R}_{j})$, represents the naive portfolio diversification rule (the 1/N portfolio) (DeMiguel et al., 2009).
@@
In the mean-variance framework, the key plug-in component is the covariance matrix, which is based exclusively on the input data set $\boldsymbol{x}$ and does not provide a guarantee of achieving portfolio diversification outside the sample. In contrast, the s-RBFN framework incorporates portfolio diversification directly into the modeling process by treating assets as hypotheses and the portfolio as an ensemble predictor, leveraging the diversity among individual predictors during training. The structured data set $\boldsymbol{R}(\varepsilon)$, constructed from predictions generated by models using the same input data as the mean-variance model, enables the use of a continuous diversification parameter $10<\varepsilon<$ 1. This allows for the creation of multiple diversified plug-in data sets from the same input and for portfolio diversification to be adjusted prior to weight optimization.
@@
Both approaches involve matrix multiplication as a core computational operation in portfolio optimization, rendering them computationally equivalent in this respect. In the mean-variance framework, least squares optimization can be applied to the input matrix $\boldsymbol{X}$ to determine portfolio weights. In the s-RBFN framework, this input matrix is replaced with the diversified prediction matrix $\boldsymbol{R}(\varepsilon)$, which is augmented through the use of basis functions. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
@@
Figure 2: Systems representation comparison of the Mean-Variance approach to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, denoted as $\boldsymbol{x}=\{\boldsymbol{x_{j}}\}_{j=1}^{M}$, and the target portfolio returns, denoted as $\boldsymbol{y}$, are assumed to be normally distributed. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{R}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x_{j}})\}_{j=1}^{M}$, parameterized by $\Theta=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, and the s-RBFN model basis functions $\boldsymbol{\Phi}$.
@@
The mean-variance approach is a fully supervised method, where the portfolio optimization process is guided by a predefined target return series. In contrast, the s-RBFN method operates within a semi-supervised learning paradigm, as it does not rely on a predefined target. Instead, the optimal portfolio weights are determined based on the hypotheses generated by the predictors and their mixtures during training. Consequently, to provide a canonical reference for the hypothesis space, it is necessary to employ an equal-weighted portfolio as a target portfolio ($\boldsymbol{y}=\frac{1}{||\boldsymbol{x}||}\boldsymbol{x}$ in Figure 2). This ensures that the decision-making is governed solely by the structured data set $\boldsymbol{R}(\varepsilon)$, thereby avoiding any dependence on a specified target.

### v2

@@ 3.1 Preliminary
Let $\boldsymbol{\hat{w}}$ be an estimator of $\boldsymbol{w}\in\mathcal{R}^{M}$ based on an i.i.d. sample from a distribution $\mathcal{D}$. The prediction error can be decomposed into bias and variance components. In this case, the bias is minimized subject to the portfolio return being above the target value, and the variance is minimized. The optimization problem can be formulated as follows.
@@
where $\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-R_{T}\right)^{2}$ represents the bias of the expected return of the portfolio compared to the target return, and $\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ represents the variance of the portfolio. $\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq R_{T}$ ensures that the expected return of the portfolio is above the target return. This setup allows to find the optimal weights $\boldsymbol{\hat{w}}$ that minimize the total prediction error of the portfolio, considering both bias and variance, while ensuring that the expected return of the portfolio is above the target value. The expression in (1) generalizes the standard MSE (when $\boldsymbol{A}=\boldsymbol{I}$), reflecting the bias-variance trade-off (Cai et al., 2024):
$MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}})=\mathcal{L}_{\boldsymbol{A}}(\mathbb{E}[\boldsymbol{\hat{w}}]-\boldsymbol{w})+\text{tr}(\boldsymbol{A}\text{Var}[\boldsymbol{\hat{w}}])$ (2)
$\boldsymbol{A}\in\mathcal{R}^{M\times M}$ is a constant positive-definite scaling matrix, $\mathcal{L}_{\boldsymbol{A}}(\boldsymbol{x})=\boldsymbol{x}^{T}\boldsymbol{A}\boldsymbol{x}$ is the $\boldsymbol{A}-norm$ of the vector $\boldsymbol{x}$, and the expectation is evaluated under the distribution $\mathcal{D}$. The first term in (2) vanishes for an unbiased estimator, and the second term vanishes if the estimator has zero variance. Now, consider $\boldsymbol{w}$ as the solution to a decision problem:
@@
where the expectation is taken under $D$. For a given $\boldsymbol{\hat{w}}$ and matrix $\boldsymbol{A}$, establish a relationship between $MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}})$ and $D_{P}(\boldsymbol{\hat{w}})$, expressed as a function $f$:
$MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}})=f(D_{P}(\boldsymbol{\hat{w}}))$ (5)
For a parameterized estimator $\boldsymbol{\hat{w}}(\delta)$, compare $\delta_{MSE}=\arg\min_{\delta}MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}}(\delta))$ with $\delta_{D}=\arg\min_{\delta}D_{P}(\boldsymbol{\hat{w}}(\delta))$ and assess the decision loss increase:
@@
3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
In this portfolio structured ensemble model (PSEM) framework, the goal is to make portfolio allocation decisions at time $t$ for an investment horizon $\Delta t=T-t$. The daily time-series returns of a portfolio for the out-of-sample holding period $\Delta t$ are denoted as $\boldsymbol{R}_{T}=\{R_{Ti}\}_{i=t}^{T}$. To approximate the expected return, a structured ensemble model predicts the portfolio time-series returns as $\boldsymbol{\hat{R}}_{T}$. This model consists of $M$ individual predictors, $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, parameterized by $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, where each predictor focuses on forecasting the return of the $j$-th component of the portfolio $\{\boldsymbol{R}_{j}\}_{j=1}^{M}$. The predictors are trained using stochastic gradient descent (SGD). During each training iteration $i$, for a given input $\{\boldsymbol{x}_{ij}\}_{i=1,j=1}^{N,M}$ and its corresponding target $R_{ij}$, the parameters of the $j$-th predictor are updated as:
where $\mathcal{F}$ is a set of specific functions of the ensemble model, $\boldsymbol{w}=\{w_{j}\}_{j=1}^{M}$ represents the optimal parameters of the ensemble, which, when properly scaled, are equivalent to portfolio weights. $\boldsymbol{\hat{R}}_{t}$ represents the predicted time series returns for the portfolio in the training data. If $\boldsymbol{\hat{R}}(\varepsilon)$ is computed from the data in period $\Delta t$, the solution is $\boldsymbol{\hat{R}}_{T}$. By incorporating the diversity parameter $\varepsilon$, the ensemble achieves a balance between prediction accuracy and diversity, enhancing its generalization capabilities in portfolio forecasting tasks. Interestingly, this diversity aligns with the concept of portfolio diversification, as originally proposed in Modern Portfolio Theory (Markowitz, 1952).
Figure 1: Structured data set formation, including diversity ($\varepsilon$), and analytical ensemble training of the s-RBFN model, described in Section 3.3 (Dominguez et al., 2025). $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{i})$ is the $j^{th}$ shooter (predictor) with input data $\boldsymbol{x}_{i}$ (for example wind, distance, etc.) and $\boldsymbol{\theta}_{j}$ her shooting skills, $\hat{R}_{ij}$ is the $i^{th}$ shoot of $j^{th}$ shooter and $\boldsymbol{\hat{R}}=\boldsymbol{\hat{R}}(\varepsilon)$ is the full set of shoots after training, $\boldsymbol{R}_{t}$ is the portfolio target for the training set (equal-weighted portfolio), and $\boldsymbol{w}$ is the s-RBFN model parameters optimized by least-squares.
@@ 3.2.1 Diversity in Ensemble Learning
where $\overset{\circ}{\boldsymbol{R}}_{j}=\arg\min\mathbb{E}_{D}[\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{R}_{j})]$ and the formula for the estimated target portfolio is in this case the arithmetic combiner,
$\boldsymbol{\hat{R}}_{t}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{R}_{j})$, the equal-weighted-portfolio ($1/N$-portfolio rule (DeMiguel et al., 2009)). This decomposition separates the expected ensemble risk into: the noise, which represents the irreducible error inherent in the problem; the average bias of the portfolio returns predictions $\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R}_{t},\overset{\circ}{\boldsymbol{R}}_{j})$; the average variance of the individual models’ predictions in the portfolio ensemble $\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}[\mathcal{L}(\overset{\circ}{\boldsymbol{R}}_{j},\boldsymbol{R}_{j})]$; the expected ambiguity term, considered as the ensemble diversity, which has the opposite sign to the other terms, $\mathbb{E}\left[\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{\hat{R}}_{t},\boldsymbol{R}_{j})\right]$, capturing the disagreement among the portfolio ensemble members in their predictions.
@@ 3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
The optimal s-RBFN parameters are obtained by least-squares with the regularization parameter $\lambda_{s}$ and target portfolio $\boldsymbol{R}_{t}$:
$\boldsymbol{w}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{R}_{t}$ (13)
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, including ensuring they sum to one and remain within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\mathbf{w}=1$ and $0\leq\mathbf{w}\leq 1$.
@@
$\boldsymbol{w}=\boldsymbol{P}\boldsymbol{R}_{t}=\boldsymbol{P}\boldsymbol{w}^{\top}\boldsymbol{x}$ (14)
with $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\top}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(m\times m)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\top}$. The target portfolio time-series returns in the training set in matrix form is given by:
$\boldsymbol{R}_{t}=\boldsymbol{w}_{\text{opt}}^{\top}\boldsymbol{x}$ (15)
with $\boldsymbol{w}_{\text{opt}}$ being the optimal weights and $\boldsymbol{x}$ the time series of the portfolio constituents. For $\boldsymbol{w}$ to be optimal in equation (13), $\boldsymbol{R}_{t}$ must be given by (15), yielding equation (14).
To ensure that the optimization is based on the shape of the constituents’ hypothesis space and the degree of diversification in the structured data set of predictions that have already been incorporated into $\boldsymbol{P}$, a canonical basis of hypotheses, each of which corresponds to one of the portfolio constituents, is selected. For that, the weights in equation (14) are selected such that the portfolio $\boldsymbol{R}_{t}$ is equal-weighted, $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the optimization and final solution for the portfolio weights are based on the information in $\boldsymbol{P}$ and not other pre-specified targets:
$\boldsymbol{w}=\boldsymbol{P}\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ (16)
The Centroid Combiner Rule, adapted to the problem of portfolio optimization, can be derived from the Bias-Variance-Diversity decomposition in (10) (Wood et al., 2024):
$\boldsymbol{\hat{R}}_{t}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{R}_{j}),$
which is the equal-weighted portfolio, described in the literature as the 1/N rule, and represents the naive portfolio diversification rule (DeMiguel et al., 2009).
@@
In the mean-variance framework, the key plug-in component is the covariance matrix, which is based exclusively on the input data set $\boldsymbol{x}$ and does not provide a guarantee of achieving portfolio diversification outside the sample. In contrast, the PSEM framework incorporates portfolio diversification directly into the modeling process by treating assets as hypotheses and the portfolio as an ensemble predictor, leveraging the diversity among individual predictors during training. The structured data set $\boldsymbol{R}(\varepsilon)$, constructed from predictions generated by models using the same input data as the mean-variance model, enables the use of a continuous diversification parameter $10<\varepsilon<$ 1. This allows for the creation of multiple diversified plug-in data sets from the same input and for portfolio diversification to be adjusted prior to weight optimization.
@@
Both approaches involve matrix multiplication as a core computational operation in portfolio optimization, rendering them computationally equivalent in this respect. In the PSEM framework using the s-RBFN model, the input data $\boldsymbol{x}$ is replaced with the diversified structured data set of predictions $\boldsymbol{R}(\varepsilon)$, which is augmented through the use of basis functions. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
@@
Figure 2: Systems representation comparison of the Mean-Variance approach to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, denoted as $\boldsymbol{x}=\{\boldsymbol{x_{j}}\}_{j=1}^{M}$, and the target portfolio returns, denoted as $\boldsymbol{y}$ in the left case are assumed to be normally distributed. The target portfolio returns in the right case are $\boldsymbol{R}_{t}$. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{R}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x_{j}})\}_{j=1}^{M}$, parameterized by $\Theta=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, and the s-RBFN model basis functions $\boldsymbol{\Phi}$.
@@
The mean-variance approach is a fully supervised method, where the portfolio optimization process is guided by a predefined target return series. In contrast, the PSEM framework operates within a semi-supervised learning paradigm, as it does not rely on a predefined target. Instead, the optimal portfolio weights are determined based on the hypotheses generated by the predictors and their mixtures during training. Consequently, to provide a canonical reference for the hypothesis space, it is necessary to employ an equal-weighted portfolio as a target portfolio ($\boldsymbol{R}_{t}=\frac{1}{||\boldsymbol{x}||}\boldsymbol{x}$ in Figure 2). This ensures that the decision-making is governed solely by the structured data set $\boldsymbol{R}(\varepsilon)$ and the s-RBFN model, thereby avoiding any dependence on a specified target.